# %First␣names␣are␣abbreviated␣in␣the␣running␣head.%If␣there␣are␣more␣than␣two␣authors,␣’et␣al.’␣is␣used.http://www.docugami.com

### v2

@@ 3.1 Motivations
The underlying assumption is that, even with a lacking baseline generating training data, weak supervision provides sufficient control over each tool to enhance its performance. To answer this, our framework takes inspiration from [8, 14, 16] and employs a planner with specialized tools to generate multi-step reasoning trajectories. By decomposing complex problems into subtasks, each of them is then solved locally with fine-tuned SLMs. This approach is particularly relevant when handling tabular contexts from business documents, where privacy concerns preclude using closed-source LLMs. Unlike other existing fine-tuned tool-augmented LLM approaches like [14, 8], MATATA uses fine-tuned specialized models without relying on GPT-4 to generate training multi-step reasoning trajectories. Prompt engineering efforts are also limited to minimize human involvement and demonstrate the system’s ability to improve and evolve with minimal initial guidance, contrary to [16, 22, 4].
@@ 4 Experiments
Table 2: Accuracy (acc.) or Exact-Match (EM) scores across evaluation datasets. Prompt Engineering (PE); Fine-Tuning (FT). Best overall scores appear underlined, best section scores appear in bold.
@@
EEDP [22] ChatGPT PE 61.88 79.73
GPT-4 PE 76.05 88.67
PoT [4] SC-Codex PE 68.1 70.2 81.8
@@
Husky[14] Llama2-7B FT 20.8 42.2 77.6
Llama3-8B FT 20.9 42.3 76.6
@@
Llama2-13B FT 71.93 77.51
Llama2-70B FT 76.81 81.42
MATATA (Ours) Phi3-mini 3.8B FT 70.10 74.44 96.66
Ministral-8B FT 77.59 77.81 98.12
4.1 Training Details
@@
Train sets are used to generate reasoning trajectories and train the tools. The best model is selected with respect to scores on the validation sets. The results on the test sets are reported in Table 2 using the relative metric for each dataset, either Accuracy or Exact Match. All datasets and splits are used entirely, as in their respective papers.
@@ 4.2 Experimental Results
Table 3: Evolution of MATATA-3.8B and MATATA-8B accuracy or exact match over fine-tuning stages; with one vs. all training set(s).
@@
Phi3-mini 3.8B baseline N/A 47.06 43.05 83.87
     + weak-sup SFT One 62.00 68.19 92.89
All 67.74 71.56 96.36
     + weak-sup KTO One 69.75 71.74 94.05
All 70.10 74.44 96.66
Ministral 8B baseline N/A 57.11 51.29 80.89
     + weak-sup SFT One 70.97 75.59 89.85
All 75.68 76.19 96.64
     + weak-sup KTO One 74.63 76.36 93.38
All 77.59 77.81 98.12
@@
On TAT-QA, MATATA-8B ranges between TAT-LLM-13B and TAT-LLM-70B, while TAT-LLM uses human-expert annotations for training. Notably, the gap between MATATA-8B and TAT-LLM-70B is only 3.61%, despite the decrease in model size. On FinQA, MATATA-8B is above all methods, and MATATA-3.8B ranges between TAT-LLM-7B and 13B, and between ChatGPT and GPT-4 with EEDP-prompting, while being much smaller than those models. On TabMWP, MATATA-8B exceeds all methods not using GPT-4, ranking 2nd on the TabMWP leaderboard, with only a 0.66% difference behind Chameleon which relies on the massively larger GPT-4. For its range of parameters and size, MATATA competes with models fine-tuned on higher-quality reasoning data, larger, or extensively prompted. This also confirms MATATA’s effectiveness, intuiting that fine-tuning a larger model with weak supervision could further improve those scores.

### v3

@@ 3.1 Motivations
The underlying assumption is that, even with a lacking baseline generating training data, weak supervision provides sufficient control over each tool to enhance its performance. To answer this, MATATA takes inspiration from [8, 14, 16] and employs a planner with specialized tools to generate multi-step reasoning trajectories. By decomposing complex problems into subtasks, each of them is then solved locally with fine-tuned SLMs. This approach is particularly relevant when handling tabular contexts from business documents, where privacy concerns preclude using closed-source LLMs. Unlike other existing fine-tuned tool-augmented LLM approaches like [14, 8], MATATA uses fine-tuned specialized models without relying on GPT-4 to generate training multi-step reasoning trajectories. Prompt engineering efforts are also limited to minimize human involvement and demonstrate the system’s ability to improve and evolve with minimal initial guidance, contrary to [16, 22, 4].
@@ 4 Experiments
4.1 Training Details
Table 2: Accuracy (acc.) or Exact-Match (EM) scores across evaluation datasets. Prompt Engineering (PE); Fine-Tuning (FT). Best overall scores appear underlined, best section scores appear in bold. On TAT-QA, scores with * are reported on the golden test set (publicly available) and others on the leaderboard test set (private).
@@ 4.1 Training Details
EEDP [22] ChatGPT PE 61.88 79.73*
GPT-4 PE 76.05 88.67*
PoT [4] SC-Codex PE 68.1 70.2* 81.8
@@
Husky[14] Llama2-7B FT 20.8 42.2* 77.6
Llama3-8B FT 20.9 42.3* 76.6
@@
Llama2-13B FT 71.93 77.5
Llama2-70B FT 76.81 81.4
MATATA (Ours) Phi3-mini 3.8B FT 70.10 74.2 96.66
Ministral-8B FT 77.59 77.6 98.13
@@
Train sets are used to generate reasoning trajectories and train the tools. The best model is selected with respect to scores on the validation sets. The results on the test sets are reported in Table 2 using the relative metric for each dataset, either Accuracy or Exact Match. All datasets and splits are used entirely, as in their respective papers. On TAT-QA, Table 2 reports scores either from the online leaderboard test set (private), or the golden test set (publicly available). They respectively contain 1,669 and 1,663 questions, with 1,657 in common. Therefore, scores are first reported from the leaderboard when available, else from their respective papers. For comparison over training stages in Table 3, the golden test set is used.
@@ 4.2 Experimental Results
Table 3: Comparison of accuracy or exact match over fine-tuning stages for MATATA-3.8B and MATATA-8B; with one vs. all training set(s). On TAT-QA, scores with * are reported on the golden test set (publicly available).
@@
Phi3-mini 3.8B baseline N/A 47.06 43.05* 83.87
     + weak-sup SFT One 62.00 68.19* 92.89
All 67.74 71.56* 96.36
     + weak-sup KTO One 69.75 71.74* 94.05
All 70.10 74.44* 96.66
Ministral 8B baseline N/A 57.11 51.29* 80.89
     + weak-sup SFT One 70.97 75.59* 89.85
All 75.68 76.19* 96.64
     + weak-sup KTO One 74.63 76.36* 93.38
All 77.59 77.81* 98.13
@@
On TAT-QA, MATATA-8B ranges between TAT-LLM-13B and TAT-LLM-70B, while TAT-LLM uses human-expert annotations for training. Notably, the gap between MATATA-8B and TAT-LLM-70B is only 3.8%, despite the decrease in model size. On FinQA, MATATA-8B is above all methods, and MATATA-3.8B ranges between TAT-LLM-7B and 13B, and between ChatGPT and GPT-4 with EEDP-prompting, while being much smaller than those models. On TabMWP, MATATA-8B exceeds all methods not using GPT-4, ranking 2nd on the TabMWP leaderboard, with only a 0.65% difference behind Chameleon which relies on the massively larger GPT-4. For its range of parameters and size, MATATA competes with models fine-tuned on higher-quality reasoning data, larger, or extensively prompted. This also confirms MATATA’s effectiveness, intuiting that fine-tuning a larger model with weak supervision could further improve those scores.
@@ 5 Conclusion
5.0.1 Acknowledgements
This work was partially supported by Mathematics of Information Technology and Complex Systems (MITACS) Accelerate, grant IT41737.